OpenStudy (anonymous):

If P(B) = 0.5, P(A or B) = 0.3, and P(A and B) = 0.6. P(A) =

OpenStudy (inkyvoyd):

If P(B) = 0.5, P(A or B) = 0.3, and P(A and B) = 0.6. P(A) = P(B) should be less than or equal to P(A or B). Something is wrong with this problem.

OpenStudy (anonymous):

@ShawnaRae_Tsenekos Are you sure your post correct?

OpenStudy (anonymous):

well this is the problemm . it doesnt show anmore than this srry .

OpenStudy (anonymous):

If your post's correct, then apply the formula for: P ( A or B) = ....

OpenStudy (anonymous):

Can you write down the formula here?

OpenStudy (anonymous):

that is the formula and the question thats all it states

OpenStudy (anonymous):

P(AorB)=P(A)+P(B)-P(A&B) so if suppose to or=0.6 &=0.3 then 0.6=0.5+P(A)-0.3 P(A)=0.4